\int \left(e^x+2^x\right)dx
1
Expert's answer
2020-09-15T17:07:50-0400

"\int(e^x+2^x)dx=\int(e^x+(e^{ln(2)})^x)dx=\\
\int(e^x+e^{x\cdot ln(2)})dx=\int e^xdx+\int e^{x\cdot ln(2)}dx=\\
e^x+\frac{1}{ln(2)}\int e^{x\cdot ln(2)}d(x\cdot ln(2))=\\
e^x+\frac{1}{ln(2)}\cdot e^{x\cdot ln(2)}+C=\\
e^x+\frac{1}{ln(2)}\cdot(e^{ln(2)})^x+C=\\
e^x+\frac{2^x}{ln(2)}+C."

Answer: "e^x+\frac{2^x}{ln(2)}+C."